# Out-of-bag prediction balls for random forests in metric spaces

### v1

@@ 2.2 Random forests in metric spaces
for each element of $\mathcal{L}_{n}$. These weights, based on Fréchet trees, are used by RFWLCFR to estimate the conditional Fréchet function $M_{\oplus}(\boldsymbol{x},y)$ in (2) using the weighted scheme in (3). Furthermore, Qiu et al., (2024) derived pointwise consistency and convergence rates based on the theory of infinite-order $U$-statistics and $U$-processes, under assumptions that include the existence and uniqueness of $m(\boldsymbol{x})$ and its estimator.
Bulté and Sørensen, (2024) performed a comprehensive analysis of RFs for their use in Fréchet regression, and proposed the Metric Random Forest (MRF), which applies a new splitting rule that alleviates the computational burden associated with calculating Fréchet means. Using the $2$-means algorithm as split function, the CART criterion (4) requires calculating (or estimating) $2\ell$ Fréchet means per split (where $\ell$ is the number of variables considered at each split), yielding $O(n\ell)$ Fréchet mean computations for the whole tree. Since computing a single Fréchet mean can be costly, particularly with non-Euclidean data, Bulté and Sørensen, (2024) suggested replacing the empirical Fréchet mean with the Fréchet medoid estimator
@@
where $S\subset[0,1]^{p}$ and $\mathcal{Y}_{n}=\{Y_{1},\ldots,Y_{n}\}$. In other words, Bulté and Sørensen, (2024) simplified the minimization problem needed to calculate the Fréchet mean by searching among the sample points, instead of the whole metric space. For this, obtaining the matrix of pairwise distances of $\mathcal{Y}_{n}$ is sufficient, a step that can be completed prior to model fitting. The Fréchet mean is replaced with Fréchet medoids within the CART splitting criterion (4) in MRFs, building on RFWLCFR. The medoid approximation in (8) was shown to yield a consistent estimation of the Fréchet population mean under conditions related to the consistency of $M$-estimators. These assumptions include the existence and uniqueness of $m(\boldsymbol{x})$ for every $\boldsymbol{x}$, and a compact metric space $\mathcal{Y}$.
@@ 2.3 Out-of-bag errors
  • In FRFs (Capitaine et al.,, 2024), the construction of an OOB prediction $\widehat{Y}_{(i)}$ is simple: select the trees built with resamples $\mathcal{L}_{n}^{*b}\in\mathcal{B}_{i}$ and aggregate their predictions,
@@
  • In RFWLCFR (Qiu et al.,, 2024), begin by computing the weights that are generated by the prediction of $Y_{i}$ from each tree. For every $b\notin\mathcal{B}_{i}$, set $\omega_{j}^{b}(\boldsymbol{X}_{i},\boldsymbol{\theta}_{b})=0$ for $j=1,\ldots,n$ and aggregate the weights as $\omega_{j}(\boldsymbol{X}_{i})=\frac{1}{|\mathcal{B}_{i}|}\sum_{b=1}^{B}\omega_{j}^{b}(\boldsymbol{X}_{i},\boldsymbol{\theta}_{b})$. Finally,
@@
  • In MRFs (Bulté and Sørensen,, 2024), the medoids used in the split criterion (4) of the trees $T_{b}$, $b\in\mathcal{B}_{i}$ are computed without $(X_{i},Y_{i})$ by construction. Thus, computing the weights $\omega_{j}$ as for RFWLCFR, and denoting $\mathcal{Y}_{n}^{(i)}=\mathcal{Y}_{n}\setminus\{Y_{i}\}$, we have that
@@ 3.1 Definition and coverage types
Our goal is to obtain estimates of the uncertainty in a RF prediction of the response. In other words, the goal is to develop confidence regions that cover the true value of the target variable with a specified probability. This will be achieved by generalizing the prediction confidence intervals proposed in Zhang et al., (2020) to responses in general metric spaces.
@@
Type I convergence is known as marginal coverage, which is studied by Lei et al., (2018) and Barber et al., (2021). In this scenario, the data set $\mathcal{L}_{n}$ and the new observation $(X,Y)\sim\mathbb{G}$ are selected at random. This is interesting in simulation studies where we assume a model $\mathbb{G}$ for the data generation and want to analyze OOB balls across different random samples. In convergence Type II, a given data set is fixed, and the coverage rate for any random pair $(X,Y)$ given the data is under analysis. Type III convergence is mainly interesting for theoretical purposes, particularly when we simulate data from a model $\mathbb{G}$ and our interest is in analyzing the prediction for a specific value of $X=x$ across different samples $\mathcal{L}_{n}$. This is the convergence type studied in Zhou and Müller, (2025) and Meinshausen, (2006). Type IV coverage is the case that is commonly found in data analysis: a training set $\mathcal{L}_{n}$ is given and the coverage rate of the response for a given value of the predictors $X=x$ is under study.
@@ 3.2 Assumptions
Continuity of $F_{R}$ in condition (c.2)(c.2.3) is required to build a probabilistic version of Polya’s Theorem in Lemma A.4, and simplifies the proofs of Lemmas A.1, A.2, and A.3 (see Section A in the SM). Finally, assumption (c.3) provides consistency of the estimators of the conditional Fréchet mean $m$. This is central in the analysis of the coverage rates, since (c.3)(c.3.1) and (c.3)(c.3.2) entail results of consistency for the RF and OOB residuals, respectively (see the proofs of Lemmas A.1 and A.2). Due to the identical distributions of $\widehat{m}_{(1)}\left(X_{1}\right),\ldots,\widehat{m}_{(n)}\left(X_{n}\right)$ by condition (c.1), assumption (c.3)(c.3.2) entails consistency of each $\widehat{m}_{(i)}\left(X_{i}\right)$ for $i=1,\ldots,n$. In fact, the consistency of $\widehat{m}$ is implied by the consistency of $\widehat{m}_{(1)}$, as it comprises a subset of the trees in the larger forest represented by $\widehat{m}$. Thus, condition (c.3)(c.3.1) is included for clarity, although it is implied by (c.3)(c.3.2).
Assumptions (c.2)(c.2.1) and (c.2)(c.2.2) adapt additive error models for Euclidean data. The traditional regression function in our framework is the conditional Fréchet mean function, and the error $R$ is just implicitly defined through condition (c.2)(c.2.2). This is important in the construction of data generation processes satisfying the above conditions, since the introduction of random noise must be done “symmetrically about m(x)m(x)” so that the regression function $m(x)$ satisfies (11) for $\mathbb{G}$-a.s. $x\in\mathcal{X}$. Furthermore, independence of $R$ from $X$ by condition (c.2)(c.2.2) can be stronger than the typical assumption of homoscedasticity in Euclidean regression. Still, it should be satisfied by construction in most models if $X$ only affects the location of $Y$ (and not its dispersion around the mean $m(X)$). Condition (c.2)(c.2.2) is needed to derive asymptotic coverage rates of Types III and IV, where the probability is conditional on the value of $X$, and is inherited from Zhang et al., (2020).
@@ 3.3.2 Sphere and hyperboloid
Suppose $\boldsymbol{Y}\mid X=x$ follows a von Mises–Fisher (vMF) distribution, $\boldsymbol{Y}\mid X=x\sim\mathrm{vMF}(\boldsymbol{m}_{0}(x),\kappa)$, where the vMF density is $f_{\mathrm{vMF}}(\boldsymbol{y};\boldsymbol{\mu},\kappa)\propto\exp{\left\{\kappa\boldsymbol{\mu}^{\top}\boldsymbol{y}\right\}}$, $\boldsymbol{y},\boldsymbol{\mu}\in\mathbb{S}^{q}$. Then, for $\kappa>0$ and every $x$, there exists a unique Fréchet mean, which coincides with $\boldsymbol{m}_{0}(x)$ (McCormack and Hoff,, 2023, Section 4.3). Furthermore, the von Mises–Fisher distribution is rotationally symmetric about the location parameter $\boldsymbol{m}_{0}(x)$. Therefore, for every $x$, the distribution of $\boldsymbol{Y}^{\top}\boldsymbol{m}_{0}(x)$ does not depend on $\boldsymbol{m}_{0}(x)$, and hence (c.2)(c.2.2) holds. A particular regression model is
@@
Hyperbolic space is the model space for negative curvature in differential geometry (McCormack and Hoff,, 2023). Consider the Minkowski pseudo-inner product $(\boldsymbol{x},\boldsymbol{y}):=-x_{1}y_{1}+\sum_{i=2}^{q+1}x_{i}y_{i}$. The unit hyperboloid $\mathbb{H}^{q}:=\left\{\boldsymbol{x}\in\mathbb{R}^{q+1}:(\boldsymbol{x},\boldsymbol{x})=-1,\,x_{1}>0\right\}$ is closely related to $\mathbb{S}^{q}$. In this surface, distances are calculated as $d_{\mathbb{H}^{q}}(\boldsymbol{x},\boldsymbol{y})=\operatorname*{arccosh}\allowbreak\left(-(\boldsymbol{x},\boldsymbol{y})\right)$. The analogous distribution to the von Mises–Fisher on the hyperboloid $\mathbb{H}^{q}$ (denoted by HvMF) has density $f_{\mathrm{HvMF}}(\boldsymbol{y};\boldsymbol{\mu},\kappa)\propto\allowbreak\exp\left\{\kappa(\boldsymbol{y},\boldsymbol{\mu})\right\}$, for $\boldsymbol{y},\boldsymbol{\mu}\in\mathbb{H}^{q}$, and $\kappa>0$, with respect to the Lebesgue measure on $\mathbb{H}^{q}$ (Barndorff-Nielsen,, 1978; Jensen,, 1981). Section D in the SM details the explicit connection of the HvMF with the vMF density. McCormack and Hoff, (2023, Section 4.3) showed that the Fréchet mean of the distribution $\mathrm{HvMF}(\boldsymbol{\mu},\kappa)$ is $\boldsymbol{\mu}$. Hence, a regression model similar to that on $\mathbb{S}^{q}$ is $\boldsymbol{Y}\mid X=x\sim\mathrm{HvMF}(\boldsymbol{m}_{0}(x),\kappa)$. For example,
@@ 3.3.3 Positive definite matrices
Consider $\mathcal{S}_{q}^{+}$, the space of Symmetric Positive Definite (SPD) matrices with real entries of size $q\times q$. The problem of equipping $\mathcal{S}_{q}^{+}$ with a proper (and computationally efficient) metric has drawn research interest in recent years (Minh and Murino,, 2018). A standard choice in statistical applications is the Affine-Invariant (AI) distance (Pennec et al.,, 2006)
@@
where $\|\cdot\|_{\mathrm{F}}$ denotes the Frobenius norm, $\log(\cdot)$ stands for the matrix logarithm, and the $\lambda_{i}$’s are the eigenvalues of $\boldsymbol{S}_{1}^{-1}\boldsymbol{S}_{2}$. This metric endows $\mathcal{S}_{q}^{+}$ with a structure of a Riemannian symmetric space, invariant under affine transformations of the variables (see, e.g., Thanwerdas and Pennec,, 2019). Expressing $d_{\mathrm{AI}}$ in terms of the eigenvalues of $\boldsymbol{S}_{1}^{-1}\boldsymbol{S}_{2}$ enables faster computation. McCormack and Hoff, (2023) gave closed formulae for the Fréchet mean of invariant families of distributions on $\mathcal{S}_{q}^{+}$. One example is the Wishart distribution. Suppose $\boldsymbol{S}\sim\mathrm{Wishart}_{q}(d,\boldsymbol{\Sigma})$, for $\boldsymbol{\Sigma}\in\mathcal{S}_{q}^{+}$. Then the Fréchet mean is
@@
Another possible metric is the Log-Cholesky (LC) distance, introduced by Lin, (2019). Given two matrices $\boldsymbol{S}_{1}$, $\boldsymbol{S}_{2}\in\mathcal{S}_{q}^{+}$, let $\boldsymbol{R}_{1}$, $\boldsymbol{R}_{2}$ denote their respective Cholesky factors. Let also $\lfloor\boldsymbol{S}\rfloor$ represent the strictly lower triangular matrix of $\boldsymbol{S}$, $\mathrm{diag}(\boldsymbol{S})$ the vector with the diagonal of $\boldsymbol{S}$, and $\log(\cdot)$ be applied entrywise. The LC distance between $\boldsymbol{S}_{1}$ and $\boldsymbol{S}_{2}$ is
@@
for $j<i$, which follows from the expectations of a $\chi_{d-j+1}$ distribution and of the logarithm of a $\chi^{2}_{d-i+1}$ distribution (Pav,, 2015). Defining $\phi:\mathcal{S}_{q}^{+}\to\mathbb{R}^{q(q+1)/2}$ as $\phi(\boldsymbol{S})=(\boldsymbol{x},\boldsymbol{y})$, the Fréchet mean under the LC distance is given by
@@
This metric is an approximation of $d_{\mathrm{AI}}(\boldsymbol{S}_{1},\boldsymbol{S}_{2})$ when $\boldsymbol{S}_{1}$ and $\boldsymbol{S}_{2}$ are sufficiently close to $\boldsymbol{I}_{q}$ (see, e.g., Theorem 2.11 in Section 2.5 of Minh and Murino, (2018)).
@@
Under the AI distance, the population prediction ball at $X=x$ is formed by the matrices $\boldsymbol{S}\in\mathcal{S}_{q}^{+}$ such that $\det(\boldsymbol{S})<\det(\boldsymbol{M}_{0}(x))e^{R_{1-\alpha}}$. To determine the ball radius, the distribution of the determinant of a Wishart random matrix (Goodman,, 1963) shows that $R_{1-\alpha}$ is the logarithm of the $(1-\alpha)$-quantile of $c_{d,q}^{-q}\prod_{i=1}^{q}W_{i}$, where $W_{i}\sim\chi^{2}_{d-i+1}$. Section C.2 in the SM illustrates the population OOB balls in $\boldsymbol{S}\in\mathcal{S}_{q}^{+}$ for different distances. Observe that the AI and LE distances provide sensible OOB balls localized at the shape of the central matrix.

### v2

@@ 2.2 Random forests in metric spaces
for each element of $\mathcal{L}_{n}$. These weights, based on Fréchet trees, are used by RFWLCFR to estimate the conditional Fréchet function $M_{\oplus}(\boldsymbol{x},y)$ in (2) using the weighted scheme in (3). Furthermore, Qiu et al. (2024) derived pointwise consistency and convergence rates based on the theory of infinite-order $U$-statistics and $U$-processes, under assumptions that include the existence and uniqueness of $m(\boldsymbol{x})$ and its estimator.
Bulté and Sørensen (2024) performed a comprehensive analysis of RFs for their use in Fréchet regression, and proposed the Metric Random Forest (MRF), which applies a new splitting rule that alleviates the computational burden associated with calculating Fréchet means. Using the $2$-means algorithm as split function, the CART criterion (4) requires calculating (or estimating) $2\ell$ Fréchet means per split (where $\ell$ is the number of variables considered at each split), yielding $O(n\ell)$ Fréchet mean computations for the whole tree. Since computing a single Fréchet mean can be costly, particularly with non-Euclidean data, Bulté and Sørensen (2024) suggested replacing the empirical Fréchet mean with the Fréchet medoid estimator
@@
where $S\subset[0,1]^{p}$ and $\mathcal{Y}_{n}=\{Y_{1},\ldots,Y_{n}\}$. In other words, Bulté and Sørensen (2024) simplified the minimization problem needed to calculate the Fréchet mean by searching among the sample points, instead of the whole metric space. For this, obtaining the matrix of pairwise distances of $\mathcal{Y}_{n}$ is sufficient, a step that can be completed prior to model fitting. The Fréchet mean is replaced with Fréchet medoids within the CART splitting criterion (4) in MRFs, building on RFWLCFR. The medoid approximation in (8) was shown to yield a consistent estimation of the Fréchet population mean under conditions related to the consistency of $M$-estimators. These assumptions include the existence and uniqueness of $m(\boldsymbol{x})$ for every $\boldsymbol{x}$, and a compact metric space $\mathcal{Y}$.
@@ 2.3 Out-of-bag errors
  • In FRFs (Capitaine et al., 2024), the construction of an OOB prediction $\widehat{Y}_{(i)}$ is simple: select the trees built with resamples $\mathcal{L}_{n}^{*b}\in\mathcal{B}_{i}$ and aggregate their predictions,
@@
  • In RFWLCFR (Qiu et al., 2024), begin by computing the weights that are generated by the prediction of $Y_{i}$ from each tree. For every $b\notin\mathcal{B}_{i}$, set $\omega_{j}^{b}(\boldsymbol{X}_{i},\boldsymbol{\theta}_{b})=0$ for $j=1,\ldots,n$ and aggregate the weights as $\omega_{j}(\boldsymbol{X}_{i})=\frac{1}{|\mathcal{B}_{i}|}\sum_{b=1}^{B}\omega_{j}^{b}(\boldsymbol{X}_{i},\boldsymbol{\theta}_{b})$. Finally,
@@
  • In MRFs (Bulté and Sørensen, 2024), the medoids used in the split criterion (4) of the trees $T_{b}$, $b\in\mathcal{B}_{i}$ are computed without $(X_{i},Y_{i})$ by construction. Thus, computing the weights $\omega_{j}$ as for RFWLCFR, and denoting $\mathcal{Y}_{n}^{(i)}=\mathcal{Y}_{n}\setminus\{Y_{i}\}$, we have that
@@ 3.1 Definition and coverage types
Our goal is to obtain estimates of the uncertainty in a RF prediction of the response. In other words, the goal is to develop confidence regions that cover the true value of the target variable with a specified probability. This will be achieved by generalizing the prediction confidence intervals proposed in Zhang et al. (2020) to responses in general metric spaces.
@@
Type I convergence is known as marginal coverage, which is studied by Lei et al. (2018) and Barber et al. (2021). In this scenario, the data set $\mathcal{L}_{n}$ and the new observation $(X,Y)\sim\mathbb{G}$ are selected at random. This is interesting in simulation studies where we assume a model $\mathbb{G}$ for the data generation and want to analyze OOB balls across different random samples. In convergence Type II, a given data set is fixed, and the coverage rate for any random pair $(X,Y)$ given the data is under analysis. Type III convergence is mainly interesting for theoretical purposes, particularly when we simulate data from a model $\mathbb{G}$ and our interest is in analyzing the prediction for a specific value of $X=x$ across different samples $\mathcal{L}_{n}$. This is the convergence type studied in Zhou and Müller (2025) and Meinshausen (2006). Type IV coverage is the case that is commonly found in data analysis: a training set $\mathcal{L}_{n}$ is given and the coverage rate of the response for a given value of the predictors $X=x$ is under study.
@@ 3.2 Assumptions
Continuity of $F_{R}$ in condition (c.2)(c.2.3) is required to build a probabilistic version of Polya’s Theorem in Lemma A.4, and simplifies the proofs of Lemmas A.1, A.2, and A.3 (see Section A in the SM). Finally, assumption (c.3) provides consistency of the estimators of the conditional Fréchet mean $m$. Qiu et al. (2024) and Bulté and Sørensen (2024) provide sufficient conditions for the consistency of RFWLCFR and MRF, respectively. Condition (c.3) is central in the analysis of the coverage rates, since (c.3)(c.3.1) and (c.3)(c.3.2) entail results of consistency for the RF and OOB residuals, respectively (see the proofs of Lemmas A.1 and A.2). Assumption (c.3)(c.3.2) is related to the notion of stability of the RF, recently studied in Wang et al. (2023). Due to the identical distributions of $\widehat{m}_{(1)}\left(X_{1}\right),\ldots,\widehat{m}_{(n)}\left(X_{n}\right)$ by condition (c.1), assumption (c.3)(c.3.2) entails consistency of each $\widehat{m}_{(i)}\left(X_{i}\right)$ for $i=1,\ldots,n$. In fact, the consistency of $\widehat{m}$ is implied by the consistency of $\widehat{m}_{(1)}$, as it comprises a subset of the trees in the larger forest represented by $\widehat{m}$. Thus, condition (c.3)(c.3.1) is included for clarity, although it is implied by (c.3)(c.3.2).
Assumptions (c.2)(c.2.1) and (c.2)(c.2.2) adapt additive error models for Euclidean data. The traditional regression function in our framework is the conditional Fréchet mean function, and the error $R$ is just implicitly defined through condition (c.2)(c.2.2). This is important in the construction of data generation processes satisfying the above conditions, since the introduction of random noise must be done “symmetrically about m(x)m(x)” so that the regression function $m(x)$ satisfies (11) for $\mathbb{G}$-a.s. $x\in\mathcal{X}$. Furthermore, independence of $R$ from $X$ by condition (c.2)(c.2.2) can be stronger than the typical assumption of homoscedasticity in Euclidean regression. Still, it should be satisfied by construction in most models if $X$ only affects the location of $Y$ (and not its dispersion around the mean $m(X)$). Condition (c.2)(c.2.2) is needed to derive asymptotic coverage rates of Types III and IV, where the probability is conditional on the value of $X$, and is inherited from Zhang et al. (2020).
@@ 3.3.2 Sphere and hyperboloid
Suppose $\boldsymbol{Y}\mid X=x$ follows a von Mises–Fisher (vMF) distribution, $\boldsymbol{Y}\mid X=x\sim\mathrm{vMF}(\boldsymbol{m}_{0}(x),\kappa)$, where the vMF density is $f_{\mathrm{vMF}}(\boldsymbol{y};\boldsymbol{\mu},\kappa)\propto\exp{\left\{\kappa\boldsymbol{\mu}^{\top}\boldsymbol{y}\right\}}$, $\boldsymbol{y},\boldsymbol{\mu}\in\mathbb{S}^{q}$. Then, for $\kappa>0$ and every $x$, there exists a unique Fréchet mean, which coincides with $\boldsymbol{m}_{0}(x)$ (McCormack and Hoff, 2023, Section 4.3). Furthermore, the von Mises–Fisher distribution is rotationally symmetric about the location parameter $\boldsymbol{m}_{0}(x)$. Therefore, for every $x$, the distribution of $\boldsymbol{Y}^{\top}\boldsymbol{m}_{0}(x)$ does not depend on $\boldsymbol{m}_{0}(x)$, and hence (c.2)(c.2.2) holds. A particular regression model is
@@
Hyperbolic space is the model space for negative curvature in differential geometry (McCormack and Hoff, 2023). Consider the Minkowski pseudo-inner product $(\boldsymbol{x},\boldsymbol{y}):=-x_{1}y_{1}+\sum_{i=2}^{q+1}x_{i}y_{i}$. The unit hyperboloid $\mathbb{H}^{q}:=\left\{\boldsymbol{x}\in\mathbb{R}^{q+1}:(\boldsymbol{x},\boldsymbol{x})=-1,\,x_{1}>0\right\}$ is closely related to $\mathbb{S}^{q}$. In this surface, distances are calculated as $d_{\mathbb{H}^{q}}(\boldsymbol{x},\boldsymbol{y})=\operatorname*{arccosh}\allowbreak\left(-(\boldsymbol{x},\boldsymbol{y})\right)$. The analogous distribution to the von Mises–Fisher on the hyperboloid $\mathbb{H}^{q}$ (denoted by HvMF) has density $f_{\mathrm{HvMF}}(\boldsymbol{y};\boldsymbol{\mu},\kappa)\propto\allowbreak\exp\left\{\kappa(\boldsymbol{y},\boldsymbol{\mu})\right\}$, for $\boldsymbol{y},\boldsymbol{\mu}\in\mathbb{H}^{q}$, and $\kappa>0$, with respect to the Lebesgue measure on $\mathbb{H}^{q}$ (Barndorff-Nielsen, 1978; Jensen, 1981). Section D in the SM details the explicit connection of the HvMF with the vMF density. McCormack and Hoff (2023, Section 4.3) showed that the Fréchet mean of the distribution $\mathrm{HvMF}(\boldsymbol{\mu},\kappa)$ is $\boldsymbol{\mu}$. Hence, a regression model similar to that on $\mathbb{S}^{q}$ is $\boldsymbol{Y}\mid X=x\sim\mathrm{HvMF}(\boldsymbol{m}_{0}(x),\kappa)$. For example,
@@ 3.3.3 Positive definite matrices
Consider $\mathcal{S}_{q}^{+}$, the space of Symmetric Positive Definite (SPD) matrices with real entries of size $q\times q$. The problem of equipping $\mathcal{S}_{q}^{+}$ with a proper (and computationally efficient) metric has drawn research interest in recent years (Minh and Murino, 2018). A standard choice in statistical applications is the Affine-Invariant (AI) distance (Pennec et al., 2006)
@@
where $\|\cdot\|_{\mathrm{F}}$ denotes the Frobenius norm, $\log(\cdot)$ stands for the matrix logarithm, and the $\lambda_{i}$’s are the eigenvalues of $\boldsymbol{S}_{1}^{-1}\boldsymbol{S}_{2}$. This metric endows $\mathcal{S}_{q}^{+}$ with a structure of a Riemannian symmetric space, invariant under affine transformations of the variables (see, e.g., Thanwerdas and Pennec, 2019). Expressing $d_{\mathrm{AI}}$ in terms of the eigenvalues of $\boldsymbol{S}_{1}^{-1}\boldsymbol{S}_{2}$ enables faster computation. McCormack and Hoff (2023) gave closed formulae for the Fréchet mean of invariant families of distributions on $\mathcal{S}_{q}^{+}$. One example is the Wishart distribution. Suppose $\boldsymbol{S}\sim\mathrm{Wishart}_{q}(d,\boldsymbol{\Sigma})$, for $\boldsymbol{\Sigma}\in\mathcal{S}_{q}^{+}$. Then the Fréchet mean is
@@
Another possible metric is the Log-Cholesky (LC) distance, introduced by Lin (2019). Given two matrices $\boldsymbol{S}_{1}$, $\boldsymbol{S}_{2}\in\mathcal{S}_{q}^{+}$, let $\boldsymbol{R}_{1}$, $\boldsymbol{R}_{2}$ denote their respective Cholesky factors. Let also $\lfloor\boldsymbol{S}\rfloor$ represent the strictly lower triangular matrix of $\boldsymbol{S}$, $\mathrm{diag}(\boldsymbol{S})$ the vector with the diagonal of $\boldsymbol{S}$, and $\log(\cdot)$ be applied entrywise. The LC distance between $\boldsymbol{S}_{1}$ and $\boldsymbol{S}_{2}$ is
@@
for $j<i$, which follows from the expectations of a $\chi_{d-j+1}$ distribution and of the logarithm of a $\chi^{2}_{d-i+1}$ distribution (Pav, 2015). Defining $\phi:\mathcal{S}_{q}^{+}\to\mathbb{R}^{q(q+1)/2}$ as $\phi(\boldsymbol{S})=(\boldsymbol{x},\boldsymbol{y})$, the Fréchet mean under the LC distance is given by
@@
This metric is an approximation of $d_{\mathrm{AI}}(\boldsymbol{S}_{1},\boldsymbol{S}_{2})$ when $\boldsymbol{S}_{1}$ and $\boldsymbol{S}_{2}$ are sufficiently close to $\boldsymbol{I}_{q}$ (see, e.g., Theorem 2.11 in Section 2.5 of Minh and Murino (2018)).
@@
Section C.2 in the SM illustrates the population OOB balls in $\boldsymbol{S}\in\mathcal{S}_{q}^{+}$ for different distances. Observe that the AI and LE distances provide sensible OOB balls localized at the shape of the central matrix.